# Erik G. Larsson and Jianan Bai

### v1

@@ I Introduction
Wireless networks are expected to serve users reliably, regardless of their locations. However, in practice, network coverage is inevitably restricted by signal attenuation, shadowing, and blockage, which are important effects, especially in higher frequency bands in 5G New Radio (NR). In addition, in many environments, the low rank of the channel limits the number of streams that can be multiplexed. To compensate for these effects, one can either choose to deploy more powerful macro base stations equipped with massive-antenna technology, or to densify the network by deploying additional small cells or use cell-free schemes. Although effective, these schemes are costly and may not offer a cost-effective and sustainable solution to the challenges of coverage.
Repeaters, also known as full-duplex relays, are devices with instantaneous amplify-and-forward operation. Compared with the deployment of additional base stations, repeaters can have extremely attractive form factors, they require no backhaul or phase synchronization, and they can operate with very low power consumption. These advantages make repeaters easy and inexpensive to deploy. Repeaters have been widely deployed in 2G, 3G, and 4G networks to establish coverage especially in tunnels, and have been consistently shown to provide capacity and coverage improvements in various scenarios [1, 2, 3, 4, 5, 6, 7, 8, 9]. In the context of 5G NR and the upcoming 6G, significant efforts have been made to standardize network-controlled repeaters in 3GPP [10], which enable more versatile functionality through the use of control information.
@@ II System Model
We consider a wireless network with one source and $N$ repeaters, each with a single antenna. The repeaters are assumed to be ideal signal amplifiers with the same amplification gain $\alpha>0$ and without any induced delay. All wireless channels are assumed to be linear time-invariant (LTI). The channel impulse response (CIR) from the source to the $n$-th repeater is denoted by $g_{n}(t)$. The CIR from the $n^{\prime}$-th repeater to the $n$-th repeater is $h_{nn^{\prime}}(t)$. The channels are assumed to be reciprocal, i.e., $h_{nn^{\prime}}(t)=h_{n^{\prime}n}(t),\forall n,n^{\prime}\in[N]$, and to have no self-interference, i.e., $h_{nn}(t)=0,\forall n$. See Fig. 1 for an illustration.
@@
where “*” represents convolution. Notice that the output consists of both the signal emitted from the source and the signals echoed among the repeaters. To facilitate the analysis, we represent the system in the Laplace domain:
@@
the input-output relationship of all repeaters is given by
$\hat{{\bm{\mathrm{y}}}}=\alpha(\hat{{\bm{\mathrm{x}}}}+\hat{{\bm{\mathrm{H}}}}\hat{{\bm{\mathrm{y}}}}),$ (4)
or equivalently
@@ III Stability Analysis
III-A Maximum Amplification Gain
When all repeaters are turned off, i.e., $\alpha=0$, there is no positive feedback, and therefore the system is stable. As we continuously increase the amplification gains of the repeaters, the system starts to become unstable for some $\alpha=\alpha_{\max}>0$. This particular value of $\alpha$ is referred to as the maximum amplification gain in the sense that one should not employ an amplification gain larger than it. We aim to find $\alpha_{\max}$ for arbitrary deployments of repeaters so that their amplification gains can be properly configured without affecting system stability. Furthermore, $\alpha_{\max}$ determines the best possible coverage improvement one can expect by deploying repeaters.
@@ III-B Lower Bound By Gershgorin Disc Theorem
By Gershgorin disc theorem [11, Th. 6.1.1], the eigenvalues of $\alpha\hat{{\bm{\mathrm{H}}}}(j\omega)$ are located in the union of discs:
@@
Therefore, a sufficient condition for the non-singularity of ${\bm{\mathrm{I}}}-\alpha\hat{{\bm{\mathrm{H}}}}(j\omega)$ is that $1\notin{\mathcal{G}}(\omega)$, i.e.,
@@ IV Case 1: Two Repeaters
Consider the case of $N=2$ interacting repeaters. Due to channel reciprocity, we have
$\hat{h}_{12}=\hat{h}_{21}=\sqrt{\beta}e^{-s\tau}.$ (11)
The inter-repeater channel transfer function matrix is
@@
therefore,
$\det({\bm{\mathrm{I}}}-\alpha\hat{{\bm{\mathrm{H}}}}_{:}(j\omega))=1-\alpha^{2}\beta e^{-j2\omega\tau}.$ (13)
@@
By recursively substituting (14) and (15) into each other, we can obtain
@@
where $p(t)$ is a “decayed” impulse train
@@ V Case 2: $N$ Repeaters on A Circle
with
$k=\min\{|n-n^{\prime}|,N-|n-n^{\prime}|\}.$ (21)
@@ V-A Eigenvalues of $\alpha\hat{{\bm{\mathrm{H}}}}$
$\hat{{\bm{\mathrm{H}}}}_{\circ}=\begin{bmatrix}0&\hat{h}_{1}&\cdots&\hat{h}_{N}&\hat{h}_{N}&\cdots&\hat{h}_{1}\\ \hat{h}_{1}&0&\ddots&&\ddots&\ddots&\\ \vdots&\ddots&\ddots\\ \hat{h}_{N}\\ \hat{h}_{N}&\ddots\\ \vdots&\ddots\\ \hat{h}_{1}\end{bmatrix}.$ (22)
@@
Notice that even in this special case where the eigenvalues of $\alpha\hat{{\bm{\mathrm{H}}}}$ are obtainable in closed form as a weighted sum of the complex exponentials $\{e^{-\omega\tau_{k}}\}$, determining whether $\lambda_{n}(\alpha,\omega)-1$ has a root for an arbitrary $\alpha$ is still computationally difficult. It appears that only in the two-repeater case in Sec. IV, it is tractable to find $\alpha_{\max}$ analytically.
@@ V-B Numerical Stability Analysis
as a measure of stability, where $\Omega$ is a discrete set of angular frequencies corresponding to a 20 MHz frequency band centered around the carrier frequency with 100 Hz spacing. The result with $N=15$ repeaters is shown in Fig. 4. One can observe that the lower bound in (9) obtained by the Gershgorin disc theorem quite accurately captures the transition point at which the system starts to become unstable.
@@ V-C Coverage Extension
or equivalently,
$\alpha=\sqrt{\frac{\beta(R)-\beta(\delta+R)}{2\beta(R)\sum_{k\in[K]}\beta(L_{k})}}.$ (32)
where we use the free-space channel gain in (27).
@@
By using (32), we can (approximately) test whether a coverage extension $\delta$ is possible by comparing the required $\alpha$ with $\alpha_{G}$ in (9) and the power constraint. In Fig. 6 we show the coverage extension achieved by $N$ repeaters when they use the amplification gain $\alpha=\min\{\alpha_{G},\gamma\}$, where $\gamma\in\{60,70,80\}$ dB represents the transmit power constraint of the repeaters. The other parameters are set to be the same as in Sec. V-B. We can observe that as the number of repeaters increases, the achieved coverage extension is first restricted by the power constraint and then by the stability requirement. The optimal deployment of repeaters depends on the power constraint: when the repeaters transmit with high power, they must be sparsely deployed to avoid instability. Quantitatively, from Fig. 6, the coverage extension that can be obtained is quite small.
@@ VI Case 3: Repeaters on A Grid
as a measure of stability when the distance between two adjacent repeaters is 200 meters. The channel model and the other parameters are the same as in Sec. V-B. We can observe that the lower bound obtained from the Gershgorin disc theorem again predicts the transition point rather accurately.
@@
Figure 8: Stability test for repeaters on a grid.

### v2

@@ I Introduction
Wireless networks are expected to serve users reliably, regardless of their locations. However, in practice, network coverage is inevitably restricted by signal attenuation, shadowing, and blockage, which are important effects, especially in higher frequency bands in 5G New Radio (NR). In addition, in many environments, the low rank of the channel limits the number of streams that can be multiplexed. To compensate for these effects, one can either choose to deploy more powerful macro base stations equipped with massive-antenna technology, or to densify the network by deploying additional small cells or use cell-free schemes. Although effective, these schemes are costly and may not offer a cost-effective solution.
Wireless repeaters are devices with instantaneous amplify-and-forward operation in full-duplex. Compared with the deployment of additional base stations, repeaters can have extremely attractive form factors, they require no backhaul or phase synchronization, and they can operate with very low power consumption. These advantages make repeaters easy and inexpensive to deploy. Repeaters have been widely deployed in 2G, 3G, and 4G networks to establish coverage especially in tunnels, and have been consistently shown to provide capacity and coverage improvements in various scenarios [1, 2, 3, 4, 5, 6, 7, 8]. In the context of 5G NR and the upcoming 6G, significant efforts have been made to standardize network-controlled repeaters in 3GPP [9], which enable more versatile functionality through the use of control information.
Repeaters should not be confused with relays. A relay receives a waveform of some duration, and re-transmits it (with or without decoding). A repeater instantaneously retransmits, with a delay of a few 100 ns at most, making it act as an ordinary channel scatterer, but with amplification [10]. As long as the repeater is reciprocal, it will be transparent to the network in a TDD multiuser MIMO system [10, 11]. An important design aspect of repeaters is to keep self-interference under control. This self-interference can be a limiting factor for the repeater gain. For single-antenna repeaters, it is the antenna-interface circuit, and for dual-antenna repeaters also the inter-antenna distance, that determine the isolation between the transmit and receive paths and hence determine the maximum gain [10, 12].
@@ II System Model
We consider a wireless network with one source and $N$ repeaters, each with a single antenna. The repeaters are assumed to be ideal signal amplifiers with the same amplification gain $\alpha>0$ and without any induced delay. All wireless channels are assumed to be linear time-invariant (LTI). The channel impulse response (CIR) from the source to the $n$-th repeater is denoted by $g_{n}(t)$. The CIR from the $n^{\prime}$-th repeater to the $n$-th repeater is $h_{nn^{\prime}}(t)$. The channels are assumed to be reciprocal, i.e., $h_{nn^{\prime}}(t)=h_{n^{\prime}n}(t),\forall n,n^{\prime}\in[N]$. Also, we assume that self-interference is removed by the antenna-interface circuit or by using sufficient antenna separation, i.e., $h_{nn}(t)=0$. See Fig. 1 for an illustration.
@@
where “*” represents convolution. [^1] Note that all signals (including the signal from the source) are amplified by the repeater and hence scaled by $\alpha$ in (1). Notice that the output consists of both the signal emitted from the source and the signals echoed among the repeaters. To facilitate the analysis, we represent the system in the Laplace domain:
@@
the input-output relationship of all repeaters is given by ${\hat{{\bm{\mathrm{y}}}}=\alpha(\hat{{\bm{\mathrm{x}}}}+\hat{{\bm{\mathrm{H}}}}\hat{{\bm{\mathrm{y}}}}),}$ or equivalently
@@ III Stability Analysis
Figure 2: Block diagram corresponding to (4).
III-A Maximum Stable Amplification Gain
When all repeaters are turned off, i.e., $\alpha=0$, there is no positive feedback, and therefore the system is stable. As we continuously increase the amplification gains of the repeaters, the system starts to become unstable for some $\alpha=\alpha_{\max}>0$. This particular value of $\alpha$ is referred to as the maximum stable amplification gain in the sense that one should not employ an amplification gain larger than $\alpha_{\max}$ to keep the system stable. We aim to find $\alpha_{\max}$ for arbitrary deployments of repeaters so that their amplification gains can be properly configured without affecting system stability. Furthermore, $\alpha_{\max}$ determines the best possible coverage improvement one can expect by deploying repeaters.
@@ III-A Maximum Stable Amplification Gain
In (5), we look for the minimum positive value of $\alpha$ such that ${\bm{\mathrm{I}}}-\alpha\hat{{\bm{\mathrm{H}}}}(j\omega)$ becomes singular for some frequency $\omega$. Note that the determinant in (5) is zero precisely when there is a zero eigenvalue. Note also that the eigenvalues are complex since $\hat{{\bm{\mathrm{H}}}}(j\omega)$ is symmetric but not Hermitean. Furthermore, note that the determinant is only zero for the “critical values” of $\alpha$ corresponding to zero eigenvalues. This means that if $\alpha$ is increased without bound, the determinant will be nonzero. However, what matters for the stability is the smallest value of $\alpha$ that makes the determinant zero.
@@ III-B Lower Bound By Gershgorin Disc Theorem
By Gershgorin disc theorem [13, Th. 6.1.1], the eigenvalues of $\alpha\hat{{\bm{\mathrm{H}}}}(j\omega)$ are located in the union of discs:
@@
(Recall that we assume zero self-interference, i.e., $\hat{h}_{nn}=0,\forall n$.) Therefore, a sufficient condition for the non-singularity of ${\bm{\mathrm{I}}}-\alpha\hat{{\bm{\mathrm{H}}}}(j\omega)$ is that $1\notin{\mathcal{G}}(\omega)$, i.e.,
@@ IV Case 1: Two Repeaters
Consider the case of $N=2$ interacting repeaters. Due to channel reciprocity, we have $\hat{h}_{12}=\hat{h}_{21}=\sqrt{\beta}e^{-s\tau}.$ The inter-repeater channel transfer function matrix is
@@
therefore, $\det({\bm{\mathrm{I}}}-\alpha\hat{{\bm{\mathrm{H}}}}_{:}(j\omega))=1-\alpha^{2}\beta e^{-j2\omega\tau}.$ (11)
@@
Recursively using (12) and (13) yields
@@
where $p(t)$ is a “decaying impulse train”
@@ V Case 2: $N$ Repeaters on A Circle
with $k=\min\{|n-n^{\prime}|,N-|n-n^{\prime}|\}.$ (19)
@@ V-A Eigenvalues of $\alpha\hat{{\bm{\mathrm{H}}}}$
$\hat{{\bm{\mathrm{H}}}}_{\circ}=\begin{bmatrix}[0]0&\hat{h}_{1}&\cdots&\hat{h}_{N}&\hat{h}_{N}&\cdots&\hat{h}_{1}\\ \hat{h}_{1}&0&\ddots&&\ddots&\ddots&\\ \vdots&\ddots&\ddots\\ \hat{h}_{N}\\ \hat{h}_{N}&\ddots\\ \vdots&\ddots\\ \hat{h}_{1}\end{bmatrix}.$ (20)
@@
Notice that even in this special case where the eigenvalues of $\alpha\hat{{\bm{\mathrm{H}}}}$ are obtainable in closed form as a weighted sum of the complex exponentials $\{e^{-\omega\tau_{k}}\}$, determining whether $\lambda_{n}(\alpha,\omega)-1$ has a root for an arbitrary $\alpha$ is still computationally difficult. It appears that only in the two-repeater case in Section IV, it is tractable to find $\alpha_{\max}$ analytically.
@@ V-B Numerical Stability Analysis
as a measure of stability, where $\Omega$ is a discrete set of angular frequencies corresponding to a 20 MHz frequency band centered around the carrier frequency with 100 Hz spacing.
The result with $N=15$ repeaters is shown in Fig. 4. One can observe that the lower bound in (8) obtained by the Gershgorin disc theorem quite accurately captures the transition point at which the system starts to become unstable – hence, the bound is quite tight in this case.
@@ V-C Coverage Extension
or equivalently, $\alpha=\sqrt{\frac{\beta(R)-\beta(\delta+R)}{2\beta(R)\sum_{k\in[K]}\beta(L_{k})}},$ (30)
where we used the free-space channel gain in (25).
@@
By using (30), we can (approximately) test whether a coverage extension $\delta$ is possible by comparing the required $\alpha$ with $\alpha_{G}$ in (8) and the power constraint. In Fig. 6 we show the coverage extension achieved by $N$ repeaters when they use the amplification gain $\alpha=\min\{\alpha_{G},\gamma\}$, where $\gamma\in\{60,70,80\}$ dB represents the transmit power constraint of the repeaters. The other parameters are set to be the same as in Section V-B. We can observe that as the number of repeaters increases, the achieved coverage extension is first restricted by the power constraint and then by the stability requirement. The optimal deployment of repeaters depends on the power constraint: when the repeaters transmit with high power, they must be sparsely deployed to avoid instability. Quantitatively, from Fig. 6, the coverage extension that can be obtained is quite small.
@@ VI Case 3: Repeaters on A Grid
as a measure of stability when the distance between two adjacent repeaters is 200 meters. The channel model and the other parameters are the same as in Section V-B. We can observe that the lower bound obtained from the Gershgorin disc theorem again predicts the transition point rather accurately.
@@
Figure 8: Values of $\alpha_{G}$ with different repeater spacings.
@@ VII Conclusions
Future Directions: The analysis may be extended to consider the effects of noise injected by the repeaters. Also, the repeaters may have different amplification gains such that $\alpha$ becomes dependent on $n$, and introduce phase shifts such that $\alpha$ becomes complex-valued (or even small time-delays such that $\alpha$ becomes dependent on $\omega$), which was not considered here. Moreover, our initial case studies assumed isotropic antennas. The effects of different antenna patterns remain to be analyzed.